In a series circuit there are 3 separate resistors of 3, 8 and 5 ohms. What is the total resistance in the circuit?
1
Expert's answer
2020-03-24T09:56:51-0400

If they are connected in parallel:


"R=3+8+5=16\ \Omega"

If they are connected in series:


"\frac{1}{R'}=\frac{1}{3}+\frac{1}{8}+\frac{1}{5}"

"R'=1.519\ \Omega"